# Efficient Bayesian Computation Using Plug-and-Play Priors for Poisson Inverse Problems

### v1

@@ Abstract
This paper introduces a novel plug-and-play (PnP) Langevin sampling methodology for Bayesian inference in low-photon Poisson imaging problems, a challenging class of problems with significant applications in astronomy, medicine, and biology. PnP Langevin sampling algorithms offer a powerful framework for Bayesian image restoration, enabling accurate point estimation as well as advanced inference tasks, including uncertainty quantification and visualization analyses, and empirical Bayesian inference for automatic model parameter tuning. However, existing PnP Langevin algorithms are not well-suited for low-photon Poisson imaging due to high solution uncertainty and poor regularity properties, such as exploding gradients and non-negativity constraints. To address these challenges, we propose two strategies for extending Langevin PnP sampling to Poisson imaging models: (i) an accelerated PnP Langevin method that incorporates boundary reflections and a Poisson likelihood approximation and (ii) a mirror sampling algorithm that leverages a Riemannian geometry to handle the constraints and the poor regularity of the likelihood without approximations. The effectiveness of these approaches is demonstrated through extensive numerical experiments and comparisons with state-of-the-art methods.
@@ 1 Introduction
$p(y|x)\propto\prod_{i=1}^{m}\exp\left\{y_{i}\log(\alpha(Ax)_{i})-\alpha(Ax)_{i}-\iota_{\mathbb{R}^{+}_{0}}(x_{i})\right\},$ (1)
@@
In this paper, we study novel PnP Markov chain Monte Carlo (MCMC) techniques suitable for Poisson imaging problems. Such techniques [28, 11, 59] are valuable because they allow forms of inference that are not accessible with PnP techniques based on optimization [23] or denoising diffusion models [40]. For example, PnP MCMC techniques can be embedded within empirical Bayesian machinery to tackle semi-blind imaging problems [38] or to automatically optimize regularization parameters [64, 61]. PnP MCMC techniques are also valuable for performing uncertainty quantification analyses [28, 34]. Unfortunately, to the best of our knowledge, the PnP MCMC techniques currently available are not suitable for Poisson imaging problems due to fundamental issues related to the poor regularity properties of (1). The aim of this paper is to address this important gap in Bayesian imaging methodology by proposing two extensions of the PnP framework [28] for Poisson imaging problems.
The remainder of this article is structured as follows: Section 2 reviews key concepts in Bayesian computation relevant to Bayesian imaging with PnP priors in the context of Gaussian observation models, and explains the main challenges in applying this approach for Poisson imaging problems. Section 3 presents two methodologies that address these difficulties and extend PnP Bayesian computation to Poisson problems: (i) the reflected PnP-SKROCK, an accelerated sampling scheme that leverages a variant of the Langevin diffusion tailored for constrained domains; and (ii) the PnP Mirror Langevin algorithm, which generalises PnP-ULA to a non-Euclidean geometry where the Poisson likelihood function is Lipschitz differentiable. In Section 4, we present an extensive experimental analysis that thoroughly explores different neural network denoiser architectures and algorithmic choices for the considered Poisson problems, as well as comparisons with alternative strategies from the state-of-the-art. Conclusions and perspectives for future work are finally reported in Section 5.
@@ 2.2 Plug-and-Play priors
[…]
@@
When dealing with Poisson noise, two key challenges hinder the direct application of PnP-ULA or PPnP-ULA. First, the continuous-time Langevin process is not well-defined because of the non-negativity constraint on $x$. Second, the gradient $x\mapsto\nabla\log p(y|x)$ for (1) is not globally Lipschitz, which is key for guaranteeing the convergence of the Langevin process to $p(x|y)$. In the next section, we discuss two approaches to deal with these challenges, each leading to new PnP algorithms.
@@ 3.1.1 The reflected overdamped Langevin diffusion process
where $\kappa_{t}$ is a local time that increases only on the boundary +$\partial\mathbb{R}^{n}_{+}$. The local time enforces the non-negativity constraint, as required by (1) [45, 39], while taking $\beta>0$ ensures that $\nabla\log{p}^{\beta}(y|x)$ is Lipschitz-continuous on the positive orthant [^1] The constant $\beta>0$ can be interpreted as a constant background noise level.. Note that setting $\beta=0$ recovers the original likelihood (1), and that the bias stemming from using $\beta>0$ can be made arbitrarily small by reducing the value of $\beta$, at the expense of convergence speed (see [39, Section 4.4] for recommendations regarding setting $\beta$).
The RSDE (7) can be approximated by a standard EM scheme, leading to a constrained ULA. Two standard ways to discretize this RSDE ensuring that the samples are always in ++$\mathbb{R}^{n}_{++}$ are using a reflected [39] or a projected Euler scheme [28]. The reflected EM scheme is defined by
@@
where $\delta$ and $\xi_{k}$ remain as previously, and |.||.|| . | denotes the component-wise absolute value. The projected EM scheme is defined by
@@ 3.1.2 Reflected and Projected PnP Langevin algorithms
PnP priors encoded by neural network denoisers can become unreliable when evaluated on out-of-distribution data [28]. Therefore, to enable the reliable use of recursions (9) and (10) for sampling, we impose an additional constraint to ensure that the iterates remain within the domain of $D_{\epsilon}$, which we henceforth denote by $C\subset\mathbb{R}^{d}$. Again, we can incorporate this constraint through reflection or projection onto $C$. This leads to two variants of PnP-ULA, namely PnP-ULA with reflection (RPnP-ULA, see Algorithm 1), and PnP-ULA with projection (PPnP-ULA, see Algorithm 2). For each algorithm we need to specify the number of iterations $N$, the noisy and blurry observation $y$, the photon level $\alpha$ (which appears within $\nabla\log p^{\beta}$), a noise level $\epsilon$ for the used denoiser, a hyper-parameter $\rho>0$, a step size $\delta$, and the set $C$, which for computational tractability we assume to be convex and compact. The choice of the step size $\delta$ for RPnP-ULA is governed by bounds given in [28], which requires $\delta<\frac{1}{3}\operatorname{Lip}(\nabla\log p_{\beta,\epsilon}(x|y))^{-1}$ where $\operatorname{Lip}(.)$ denotes the Lipschitz constant. PPnP-ULA admits a larger step size, at the expense of additional estimation bias [28].
@@
Compute $l_{s}=(s-0.5)^{2}(2-4/3\eta)-1.5\vspace{0.2cm}$
[…]
@@
[…]
@@ 3.2 A Riemannian overdamped Langevin diffusion process
For Poisson inverse problems, a natural choice for $\phi$ is the Burg’s entropy (or maximum entropy) [6], given by
$\phi(x)=\sum_{i=1}^{n}\log(x_{i}),$ (12)
as the likelihood (1) is smooth w.r.t. to this choice of $\phi$ (see, e.g.,  [3] for details). Again, there are many ways to obtain a discrete sampling scheme from a mirror Langevin SDE [33, 27, 68, 1, 22, 19]. We choose the EM scheme studied in [68]
${X}_{k+1}:=\nabla\phi^{*}\left(\nabla\phi({X}_{k})-\delta\nabla\log\pi({X}_{k})+\sqrt{2\delta[\nabla^{2}\phi({X}_{k})]}\xi_{k+1}\right)\,.$ (13)
@@ 4.1 Choice of PnP prior for Poisson imaging problems
We also consider B-DRUNet [23], a Bregman generalization of the Euclidean proximal gradient step denoiser [25]. This denoiser is based a generalised form of Tweedie’s identity, suitable for denoising problems with non-Gaussian noise distributions. More precisely, given a noisy observation $z$ of an image $x$, then [23]
@@
where $\phi$ is a $\mathcal{C}^{2}$ convex potential of Legendre type, related to the distribution of the noise, and $\gamma$ a scalar representing the noise level. Note that when $\phi$ is quadratic, (14) reduces to the conventional Tweedie identity (4) for Gaussian denoising. It is useful to set $\phi$ to match the algebraic form of the likelihood function. In our case, this leads to the Burg entropy (12) introduced previously, which is related to inverse gamma noise (see [23] for details). Given a trained Bregman denoiser $B_{\gamma}(z)\approx\hat{x}_{\text{MMSE}}(z)$, we obtain the required approximation to the score prior score function as $-\nabla\log p_{\gamma}(z)\approx\gamma\nabla^{2}\phi(z)(z-{B}_{\gamma}(z))$. Note that in this case, we have $B_{\gamma}(z)=z-(\nabla\phi(z))^{-1}\cdot\nabla g_{\gamma}(z)$, which simplifies to $\nabla g_{\gamma}\approx-\frac{1}{\gamma}\nabla\log p_{\gamma}$. Thus, we can use our algorithms without any modification by setting $\epsilon=\frac{1}{\gamma}$ and $\nabla\log p_{\epsilon}(x)=-\frac{1}{\epsilon}\nabla g_{1/\epsilon}(x)$.
@@ 4.2.1 Experimental set up
With regards to step size values, for PPnP-ULA, [28] suggests setting $\delta\in(0,\delta_{L})$ with $\delta_{L}=1/(L_{p^{\beta}}+L_{\epsilon}/\epsilon)$, where $L_{p^{\beta}}=\alpha^{2}\cdot(\max(y)/\beta^{2})\cdot||AA^{T}||$ is the Lipschitz constant of ${\log p^{\beta}}$ and $L_{\epsilon}$ is the Lipschitz constant of the denoiser; LMMO and Prox-DRUNet have been trained so that $L_{\epsilon}=1$, [^9] The training loss of LMMO and Prox-DRUNet includes a penalization term on the spectral norm of the Jacobian, for more details see [25, 44]. while for GS-DRUNet $51<L_{\epsilon}<$ 5 [25]. However, while this choice is suitable for Gaussian imaging problems [28], we find this choice of step size overly conservative for Poisson problems and set $\delta=0.2\cdot 10^{2}\cdot\delta_{L}$, which in our experience provides a good accuracy-speed trade-off. To the best of our knowledge, there is no theory for setting the step size of PnP-MLA (Algorithm 4) under a PnP prior or a non-convex setting. We set the step size in PnP-MLA as $\delta=10^{-5}\simeq(2\cdot 10^{2})\cdot\delta_{L}$ by using grid search and optimizing for PSNR. We run the algorithms for $10^{6}$ iterations to stress-test the stability of the algorithms. Using grid search, the parameter $\epsilon$ of the LMMO, GS-DRUNet, Prox-DRUNet denoisers $D_{\epsilon}$ is set to $\epsilon=(20/255)^{2}$ and $\rho=1$. For the B-DRUNet denoiser $B_{\gamma}$, we set $\gamma=30$ and $\rho=2$. We set $C=\{x:0\leq x_{i}\leq 1\}$, as this corresponds to the training domain of the denoisers.

### v2

@@ Abstract
This paper studies plug-and-play (PnP) Langevin sampling strategies for Bayesian inference in low-photon Poisson imaging problems, a challenging class of problems with significant applications in astronomy, medicine, and biology. PnP Langevin sampling offers a powerful framework for Bayesian image restoration, enabling accurate point estimation as well as advanced inference tasks, including uncertainty quantification and visualization analyses, and empirical Bayesian inference for automatic model parameter tuning. Herein, we leverage and adapt recent developments in this framework to tackle challenging imaging problems involving weakly informative Poisson data. Existing PnP Langevin algorithms are not well-suited for low-photon Poisson imaging due to high solution uncertainty and poor regularity properties, such as exploding gradients and non-negativity constraints. To address these challenges, we explore two strategies for extending Langevin PnP sampling to Poisson imaging models: (i) an accelerated PnP Langevin method that incorporates boundary reflections and a Poisson likelihood approximation and (ii) a mirror sampling algorithm that leverages a Riemannian geometry to handle the constraints and the poor regularity of the likelihood without approximations. The effectiveness of these approaches is evaluated and contrasted through extensive numerical experiments and comparisons with state-of-the-art methods. The source code accompanying this paper is available at https://github.com/freyyia/pnp-langevin-poisson.
@@ 1 Introduction
$p(y|x)\propto\prod_{i=1}^{m}\exp\left\{y_{i}\log(\alpha(Ax)_{i})-\alpha(Ax)_{i}-\iota_{\mathbb{R}^{+}_{0}}(x_{i})\right\}.$ (1)
@@
In this paper, we study novel PnP Markov chain Monte Carlo (MCMC) techniques suitable for Poisson imaging problems. Such techniques [35, 14, 66] are valuable because they allow forms of inference that are not accessible with PnP techniques based on optimization [29] or denoising diffusion models [47]. For example, PnP MCMC techniques can be embedded within empirical Bayesian machinery to tackle semi-blind imaging problems [45] or to automatically optimize regularization parameters [73, 69]. PnP MCMC techniques are also valuable for performing uncertainty quantification analyses [35, 41]. Unfortunately, to the best of our knowledge, the PnP MCMC techniques currently available are not suitable for Poisson imaging problems due to fundamental issues related to the poor regularity properties of (1).
The primary goal of this paper is to make Plug-and-Play (PnP) MCMC a viable and powerful tool for Bayesian inference in challenging Poisson imaging problems, addressing an important gap where existing PnP Langevin methods are ill-suited due to the poor regularity of the likelihood and constraints. Our contributions are threefold. First, we systematically adapt and extend state-of-the-art Euclidean and non-Euclidean samplers to the specific difficulties of the PnP Poisson setting, including the first implementation and investigation of a PnP Mirror Langevin algorithm (PnP-MLA) for an imaging problem. Second, we present the first comprehensive empirical study that directly compares these different geometric and accelerated MCMC methodologies (reflected, projected, mirror) within a unified PnP framework, yielding practical insights beyond the scope of theoretical analysis. Finally, this comparative analysis provides a set of recommendations for practitioners on selecting the appropriate algorithm and denoiser for various low-photon regimes.
The remainder of this article is structured as follows: Section 2 reviews key concepts in Bayesian computation relevant to Bayesian imaging with PnP priors in the context of Gaussian observation models, and explains the main challenges in applying this approach for Poisson imaging problems. Section 3 presents two methodologies that address these difficulties and extend PnP Bayesian computation to Poisson problems: (i) the reflected PnP-SKROCK, an accelerated sampling scheme that leverages a variant of the Langevin diffusion tailored for constrained domains; and (ii) the PnP Mirror Langevin algorithm, which generalizes PnP-ULA to a non-Euclidean geometry where the Poisson likelihood function is Lipschitz differentiable. In Section 4, we present an extensive experimental analysis that thoroughly explores different neural network denoiser architectures and algorithmic choices for the considered Poisson problems, as well as comparisons with alternative strategies from the state-of-the-art. A thorough discussion, conclusions and perspectives for future work are finally reported in Sections 5 and 6, respectively.
@@ 2.2 Plug-and-Play priors
where $p_{\epsilon}(x)$ is a smooth approximation of $p(x)$ defined as $p_{\epsilon}(x)=\int k_{\epsilon}(x,\tilde{x})p(\tilde{x})d\tilde{x}$, obtained via convolution with a Gaussian smoothing kernel $k_{\epsilon}$ of bandwidth $\sqrt{\epsilon}$. As $\epsilon$ decreases, the approximation $p_{\epsilon}$ converges to $p$, at the expense of reduced smoothness.
@@
When dealing with Poisson noise, two main challenges hinder the direct application of PnP-ULA or PPnP-ULA. First, the continuous-time Langevin process is not well-defined because of the non-negativity constraint on $x$. Second, the gradient $x\mapsto\nabla\log p(y|x)$ for (1) is not globally Lipschitz, which is important for guaranteeing the convergence of the Langevin process to $p(x|y)$. In the next section, we discuss two approaches to deal with these challenges, each leading to new PnP algorithms.
@@ 3.1.1 The reflected overdamped Langevin diffusion process
where $\kappa_{t}$ is a local time that increases only on the boundary $\partial\mathbb{R}^{n}_{+}$. The local time enforces the non-negativity constraint, as required by (1) [52, 46], while taking $\beta>0$ ensures that $\nabla\log{p}^{\beta}(y|x)$ is Lipschitz-continuous on the positive orthant [^1] The constant $\beta>0$ can be interpreted as a constant background noise level.. Note that setting $\beta=0$ recovers the original likelihood (1), and that the bias stemming from using $\beta>0$ can be made arbitrarily small by reducing the value of $\beta$, at the expense of convergence speed (see [46, Section 4.4] for recommendations regarding setting $\beta$).
The RSDE (7) can be approximated by a standard EM scheme, leading to a constrained ULA. Two standard ways to discretize this RSDE ensuring that the samples are always in $\mathbb{R}^{n}_{++}$ are using a reflected [46] or a projected Euler scheme [35]. The reflected EM scheme is defined by
@@
where $\delta$ and $\xi_{k}$ remain as previously, and $|.|$ denotes the component-wise absolute value. The projected EM scheme is defined by
@@ 3.1.2 Reflected and Projected PnP Langevin algorithms
PnP priors encoded by neural network denoisers can become unreliable when evaluated on out-of-distribution data [35]. Therefore, to enable the reliable use of recursions (9) and (10) for sampling, we impose an additional constraint to ensure that the iterates remain within the domain of $D_{\epsilon}$, which we henceforth denote by $C\subset\mathbb{R}^{d}$ and assume to be convex and compact. This constraint can be enforced using either reflection on the boundary $\partial C$ or projection onto $C$, leading to two variants of the PnP-ULA. In both cases, the Langevin update step targets a posterior with log-density of the form
$\log p_{\beta,\epsilon}(x|y)\propto\log p^{\beta}(y|x)+\rho\log p_{\epsilon}(x)\,,\quad\forall x\in C$
where $\rho>0$ is a regularization parameter. The prior score, $\nabla\log p_{\epsilon}(x)$, is related to the denoiser $D_{\epsilon}$ using Tweedie’s identity (4) as $\nabla\log p_{\epsilon}(x)=\frac{1}{\epsilon}(D_{\epsilon}(x)-x)$. For specific network architectures, such as gradient step denoisers discussed in Section 4.1, this expression can be further simplified.
The resulting algorithms are PnP-ULA with reflection (RPnP-ULA, see Algorithm 1), and PnP-ULA with projection (PPnP-ULA, see Algorithm 2). These algorithms require specifying the number of iterations $N$, the observation $y$, the photon level $\alpha$ (which appears within $\nabla\log p^{\beta}$), a noise level $\epsilon$ for the used denoiser, and a step size $\delta$. For RPnP-ULA, a reflection operator $\mathcal{R}_{C}$ is applied after each update step to keep the iterates within the set $C$. The choice of the step size $\delta$ for RPnP-ULA is governed by stability bounds given in [35], which requires $\delta<\frac{1}{3}\operatorname{Lip}(\nabla\log p_{\beta,\epsilon}(x|y))^{-1}$ where $\operatorname{Lip}(.)$ denotes the Lipschitz continuity constant. In contrast, PPnP-ULA uses a projection operator $\Pi_{C}$ to enforce the constraint. PPnP-ULA admits a larger step size than its reflected counterpart, at the expense of introducing additional estimation bias [35]. For a constraint set $C\in[a,b]^{d}$, the projection operator is applied component-wise. The projection operator is given as $\Pi_{C}(x)_{i}=\min(\max(a,x_{i}),b)$, and the reflection operator can be defined in terms of the projection as $\mathcal{R}_{C}(x)_{i}=2\Pi_{C}(x)_{i}-x_{i}$.
@@
Compute $l_{s}=(s-0.5)^{2}(2-4/3\eta)-1.5\vskip 5.69046pt$
Compute $\omega_{0}=1+\dfrac{\eta}{s^{2}},\hskip 9.95863pt\omega_{1}=\dfrac{T_{s}(\omega_{0})}{T^{{}^{\prime}}_{s}(\omega_{0})},\hskip 9.95863pt\mu_{1}=\dfrac{\omega_{1}}{\omega_{0}},\hskip 9.95863pt\nu_{1}=s\omega_{1}/2,\hskip 9.95863ptk_{1}=s\omega_{1}/\omega_{0}\vskip 5.69046pt$
@@
   Compute $\mu_{j}=\dfrac{2\omega_{1}T_{j-1}(\omega_{0})}{T_{j}(\omega_{0})},\hskip 9.95863pt\nu_{j}=\dfrac{2\omega_{0}T_{j-1}(\omega_{0})}{T_{j}(\omega_{0})},\hskip 9.95863ptk_{j}=-\dfrac{T_{j-2}(\omega_{0})}{T_{j}(\omega_{0})}=1-\nu_{j}\vskip 5.69046pt$
@@ 3.2 A Riemannian overdamped Langevin diffusion process
It is important to choose a mirror map $\phi(x)$ that is well-suited to the structure of the potential $U(x)=-\log\pi(x)$. The convergence analysis in [78] is based on the notions of relative Lipschitz smoothness and relative strong convexity of $U$ with respect to $\phi$. This condition is satisfied if there exists constants $0<m\leq L<\infty$ such that $\forall x\in\mathcal{X}$
$m\nabla^{2}\phi(x)\;\preccurlyeq\;\nabla^{2}U(x)\;\preccurlyeq\;L\nabla^{2}\phi(x),$ (12)
[…]
For Poisson inverse problems, the negative log-likelihood (1) is not globally Lipschitz smooth in the Euclidean sense, however a natural choice for the mirror map $\phi$ is the Burg’s entropy (or maximum entropy) [9], given by
$\phi(x)=\sum_{i=1}^{n}\log(x_{i}).$ (13)
The Hessian of this map, given by $\nabla^{2}\phi(x)=\operatorname{diag}(1/x_{1}^{2},\dots,1/x_{n}^{2})$, strongly penalizes values near the boundary of the positive orthant. In addition, (1) is relatively smooth and strongly convex with respect to this choice of $\phi$ (see [6, 29] for details), and thus the conditions on the likelihood for convergence of the mirror Langevin method are satisfied. This relationship extends to the contractive and Lipschitz-constrained PnP priors that we consider (see Section 2.2) due to the effective Lipschitz constant.
There are many ways to obtain a discrete sampling scheme from a mirror Langevin SDE [40, 34, 78, 2, 28, 23]. We choose the EM scheme studied in [78]
${X}_{k+1}:=\nabla\phi^{*}\left(\nabla\phi({X}_{k})+\delta\nabla\log\pi({X}_{k})+\sqrt{2\delta[\nabla^{2}\phi({X}_{k})]}\xi_{k+1}\right)\,.$ (14)
@@ 4.1 Choice of PnP prior for Poisson imaging problems
There exists a main difference between the GS-DRUNet and Prox-DRUNet variants [30, 31]. While both share the same DRUNet architecture and are initially trained by minimizing a standard $L^{2}$ loss,
$\mathcal{L}(\epsilon)=\mathbb{E}_{x\sim p,\nu_{\epsilon}\sim\mathcal{N}(0,\epsilon^{2})}\left[\left\|D_{\epsilon}(x+\nu_{\epsilon})-x\right\|^{2}\right],$
with $p$ denoting the distribution of clean training images, the Prox-DRUNet proposed in [31] undergoes an additional fine-tuning stage. This fine-tuning uses a modified objective that includes a regularizer on the spectral norm $||.||_{S}$ of the Jacobian $\nabla^{2}g_{\epsilon}$,
$\mathcal{L}_{S}(\epsilon)=\mathbb{E}_{x\sim p,\nu_{\epsilon}\sim\mathcal{N}(0,\epsilon^{2})}\left[\|D_{\epsilon}(x+\nu_{\epsilon})-x\|^{2}\\ +\mu\max(\|\nabla^{2}g_{\epsilon}(x+\nu_{\epsilon})\|_{S},1-\varepsilon)\right],$
parametrized by scalars $\mu$ and $\varepsilon=0.1$. The purpose of this additional term is to explicitly enforce that the operator $\nabla g_{\epsilon}$ is contractive (i.e., has a Lipschitz constant less than 1). This requires replacing the original ReLU activations with smooth alternatives like softplus to ensure the necessary differentiability. This fine-tuning with spectral norm regularization is the primary difference between Prox-DRUNet and GS-DRUNet, and we hypothesize it contributes to the often more stable and superior performance of Prox-DRUNet.
We also consider B-DRUNet [29], a Bregman generalization of the Euclidean proximal gradient step denoiser [31]. This denoiser is based a generalized form of Tweedie’s identity, suitable for denoising problems with non-Gaussian noise distributions. More precisely, given a noisy observation $z$ of an image $x$, then [29]
@@
where $\phi$ is a $\mathcal{C}^{2}$ convex potential of Legendre type, related to the distribution of the noise, and $\gamma$ a scalar representing the noise level. Note that when $\phi$ is quadratic, (15) reduces to the conventional Tweedie identity (4) for Gaussian denoising. It is useful to set $\phi$ to match the algebraic form of the likelihood function. In our case, this leads to the Burg entropy (13) introduced previously, which is related to inverse gamma noise (see [29] for details). Given a trained Bregman denoiser $B_{\gamma}(z)\approx\hat{x}_{\text{MMSE}}(z)$, we obtain the required approximation to the score prior score function as $\nabla\log p_{\gamma}(z)\approx\gamma\nabla^{2}\phi(z)(z-{B}_{\gamma}(z))$. Note that in this case, we have $B_{\gamma}(z)=z-(\nabla\phi(z))^{-1}\cdot\nabla g_{\gamma}(z)$, which simplifies to $\nabla g_{\gamma}\approx-\frac{1}{\gamma}\nabla\log p_{\gamma}$. Thus, we can use our algorithms without any modification by setting $\epsilon=\frac{1}{\gamma}$ and $\nabla\log p_{\epsilon}(x)=-\frac{1}{\epsilon}\nabla g_{1/\epsilon}(x)$.
@@ 4.2.1 Experimental set up
An important aspect of this empirical study is the selection of step size $\delta$ for each algorithm, as theoretical findings are restrictive or unavailable for this problem class. For the Euclidean algorithms (PPnP-ULA, RPnP-SKROCK), existing theoretical bounds [35] are highly conservative for Poisson problems. As our goal is to investigate practical performance limits, we deliberately explore step sizes that exceed these bounds to find a more effective accuracy-speed trade-off. To the best of our knowledge, there is no theory for setting the step size of PnP-MLA (Algorithm 4) under a PnP prior or a non-convex setting. This remains an open research problem. Consequently, we adopted an empirical approach, selecting $\delta$ for PnP-MLA via grid search to optimize PSNR. Therefore, the step sizes used herein were chosen to enable a robust and fair empirical comparison of the algorithms’ stability and performance at their practical limits.
[…]